OpenStudy (oswaldmurphy):

Can someone help me understand exactly what a polonomial is? Please explain it, not a link please!

TheSmartOne (thesmartone):

A polynomial is made up of the sum of monomials. Monomials have 1 term. Like 8x^2y^2 You can also have a fraction in a monomial BUT as long as the denominator is a number not a variable, it is considered a monomial \(\ \frac{8x^2}{5}\) is a monomial \(\ \frac{2x^2}{y}\) is not a monomial because \(\ \frac{2x^2}{y} = 2x^2y^{-1}\) and the variable should not have a negative exponent Now, if a polynomial is made up of two monomials, it is called a binomial And if it is made up of three monomials, it is called a trinomial. If it's made up of more terms, you just call it a polynomial that has a degree of whatever the highest degree of the monomial is mhmm, does that make sense? :)

TheSmartOne (thesmartone):

I can go more in to it, and more around it I just taught this concept the other day to my students xD

OpenStudy (oswaldmurphy):

Makes perfect since. Can you go more into the degrees and such?

TheSmartOne (thesmartone):

Degree of a monomial is adding the power of each of the VARIABLES Degree of a polynomial is the the highest degree there is of one of the monomials the degree of 8x^2y is 3 x^2 and y^1 2 + 1 = 3 the degree is 3 the degree of 8x^2 + 2x + 1 8x^2 has a degree of 2 2x has a degree of 1 because x^1 and 1 has a degree of 0 because 1 * x^0 = 1 * 1 = 1 the highest degree of a monomial in the polynomial is 2, which means the degree is 2